Prashnikaप्रश्निका
Code
About
‹ Back to the paper
Mathematics20184 marksProbability distributionProbability

Answer the following about the probability distribution.

Let X denote the number of hours you study during a randomly selected school day. The probability…

Let $X$ denote the number of hours you study during a randomly selected school day. The probability that $X$ can take the values '$x$' has the following form, where '$k$' is some unknown constant. $P(X = x) = \begin{cases} 0.1, & \text{if } x = 0 \\ kx, & \text{if } x = 1 \text{ or } 2 \\ k(5-x), & \text{if } x = 3 \text{ or } 4 \\ 0, & \text{otherwise} \end{cases}$
(a)[1]
Find the value of '$k$'.
(b)[3]
What is the probability that you study: (i) at least two hours? (ii) exactly two hours? (iii) at most 2 hours?

Answer

Answer (a)

The probabilities add up to $1$: $0.1 + k + 2k + 2k + k = 1$. $0.1 + 6k = 1$ Hence $k = 0.15$.

Final answer: $0.15$

Answer (b)

(i) $P(X \ge 2) = 2k + 2k + k = 5k = 0.75$. (ii) $P(X = 2) = 2k = 0.30$. (iii) $P(X \le 2) = 0.1 + k + 2k = 0.1 + 0.45 = 0.55$. Hence the probabilities are $0.75$, $0.30$ and $0.55$.

Final answer: (i) $0.75$; (ii) $0.30$; (iii) $0.55$